OpenStudy (anonymous):

Anybody have any clue how to do this: This exercise uses the radioactive decay model. A wooden artifact from an ancient tomb contains 65% of the carbon-14 that is present in living trees. How long ago was the artifact made? (The half-life of carbon-14 is 5730 years. Round your answer to the nearest whole number.) ________yr

OpenStudy (anonymous):

yeah we can do this

OpenStudy (anonymous):

put \[.65=(\frac{1}{2})^{\frac{t}{5700}}\] and solve for t

OpenStudy (anonymous):

you get \[\frac{t}{5700}=\frac{\ln(.65)}{\ln(\frac{1}{2})}\] by the change of base formula

OpenStudy (anonymous):

so \[t=\frac{5700\times \ln(.65)}{\ln(.5)}\]

OpenStudy (anonymous):

3542?

OpenStudy (anonymous):

i put it in a calculator and get 3542 rounded

OpenStudy (anonymous):

yup!